p − q = log( x + y)
1
Expert's answer
2020-10-29T14:29:52-0400

"\displaystyle


p - q = \log(x + y)\\

\textsf{The Langrange's auxiliary equation is}\\

\frac{\mathrm{d}x}{1} = \frac{\mathrm{d}y}{-1} = \frac{\mathrm{d}z}{\log(x + y)}\\


\textsf{Choosing}\, (1, 1, 0)\, \textsf{as multipliers}\\


\mathrm{d}x + \mathrm{d}y = 0\\


\textsf{Integrating both sides}\\


\int \mathrm{d}x + \int \mathrm{d}y = C_1\\


x + y = C_1\\



\textsf{Comparing the first and the last equation}\\



\frac{\mathrm{d}x}{1} = \frac{\mathrm{d}z}{\log(x + y)}\\


\int \mathrm{d}x = \int\frac{\mathrm{d}z}{\log(x + y)}\\

\int \mathrm{d}x = \int\frac{\mathrm{d}z}{\log(C_1)}\\

x = \frac{z + C_2}{\log(C_1)}\\

x\log(C_1) = z + C_2\\


x\log(C_1) - z = C_2\\


x\log(x + y)- z = C_2, \hspace{0.5cm}\{C_1 = x + y\}\\



\textsf{Thus, the solution to the partial}\\\textsf{differential equations is}\\


\phi(x + y, x\log(x + y) - z) = 0"